# Segmenting Action-Value Functions Over Time-Scales in SARSA using TD($\Delta$)

### v1

@@ 4.4 SARSA TD($\lambda,\Delta$)
Eq.17 introduces the $\lambda$-return[23, 1], which combines rewards over multiple steps to create a target for TD($\lambda$) updates. The $\lambda$-return $G_{t}^{\gamma,\lambda}$ is defined as:
$G_{t}^{\gamma,\lambda}(s_{t},a_{t})=\hat{Q}_{\gamma}(s_{t},a_{t})+\sum_{k=0}^{\infty}(\lambda\gamma)^{k}\delta_{t+k}^{\gamma}$ (17)
@@
$\mathrm{T}_{\lambda}Q(s_{t},a_{t})=Q(s_{t},a_{t})+(I-\lambda\gamma P)^{-1}(\mathrm{T}Q(s_{t},a_{t})-Q(s_{t},a_{t}))$ (18)
Here,P is the transition matrix for state-action pairs under the policy $\pi$, and Q is the action-value function. Similarly, Eq.19 defines the $\lambda$-return specific to delta estimators in TD($\lambda,\Delta$), denoted as $G_{t}^{z,\lambda_{z}}$ for each delta estimator $W_{z}$:
$G_{t}^{z,\lambda_{z}}:=\hat{W}_{z}(s_{t},a_{t})+\sum_{k=0}^{\infty}(\lambda_{z}\gamma_{z})^{k}\delta_{t+k}^{z}$ (19)
where $\delta_{t}^{0}:=\delta_{t}^{\gamma_{0}}$ and $\delta_{t}^{z}:=(\gamma_{z}-\gamma_{z-1})\hat{Q}_{\gamma_{z-1}}(s_{t+1},a_{t+1})+\gamma_{z}\hat{W}_{z}(s_{t+1},a_{t+1})-\hat{W}_{z}(s_{t+1},a_{t+1})$ are the TD-errors.
@@ 4.5 SARSA TD($\lambda,\Delta$) with Generalized Advantage Estimation (GAE)
Generalized Advantage Estimation (GAE) [28] seeks to compute the advantage function by aggregating multi-step TD errors. To apply this approach to delta estimators $W_{z}(s_{t},a_{t})$, we compute advantage estimates $A^{\Delta}(s_{t},a_{t})$ for each time-scale. Each advantage estimate $A^{\Delta}(s_{t},a_{t})$ employs a multi-step TD error pertinent to the delta estimator $W_{z}$, expressed as follows:
$A^{\Delta}(s_{t},a_{t})=\sum_{k=0}^{T-1}(\lambda_{z}\gamma_{z})^{k}\delta_{t+k}^{\Delta}$ (20)
where $\delta_{t+k}^{\Delta}:=r_{t}+\gamma_{z}\sum_{z=0}^{Z}\hat{W}_{z}(s_{t+1},a_{t+1})-\sum_{z=0}^{Z}\hat{W}_{z}(s_{t},a_{t})$.
We utilize $\gamma_{Z}$ as the discount factor and the sum of all W estimators as a surrogate for $Q_{\gamma z}$. This objective is applicable to PPO by implementing the policy update from Eq.4 and replacing A with $A^{\Delta}$. Additionally, to train each $W_{z}$, we utilize a truncated form of their corresponding $\lambda$-return, as outlined in Eq.19. For more information, see Algorithm 2.
@@
Inputs: Pick out the discount factors ($\gamma_{0},\gamma_{1},...,\gamma_{Z}$), bootstrapping steps ($k_{0},k_{1},...,k_{Z}$), and learning rates ($\alpha_{0},\alpha_{1},...,\alpha_{Z}$).
Set the initial value of $W_{z}(s,a)=0$ for all states s, actions a, and scales z.
Initialize policy $\omega$, and values $\theta^{z}\forall z$
@@
     for t = 0, 1, 2,… do $\triangleright$ Loop for each time step
         Take action $a_{t}$, observe reward $r_{t}$ and next state $s_{t+1}$.
         Select next action $a_{t+1}$ based on a policy.
         for z = 0, 1, …, Z do
              if $t\geq T$ then
                  $G^{z,\lambda_{z}}\leftarrow\hat{W}_{z}(s_{t-T},a_{t-T})+\sum_{k=0}^{T-1}(\lambda_{z}\gamma_{z})^{k}\delta_{t-T+k}^{z}\forall z$triangleright Computing multi-step return $G^{z,\lambda_{z}}$ and TD-error $\delta_{t-T+k}^{z}$ using Eq.19.
@@
         for z = 0, 1, 2,…, Z do
              $\hat{W}_{z}(s_{t-T},a_{t-T})\leftarrow\hat{W}_{z}(s_{t-T},a_{t-T})+\alpha_{z}\left(G^{z,\lambda_{z}}-\hat{W}_{z}(s_{t-T},a_{t-T})\right)$
              $A^{\Delta}=\sum_{k=0}^{T-1}(\lambda_{z}\gamma_{z})^{k}\delta_{t-T+k}^{\Delta}$triangleright Where $A^{\Delta}$ and $\delta_{t-T+k}^{\Delta}$ are computed using Eq.20.
              $\theta^{z}\leftarrow\theta^{z}+\alpha_{z}\left(G^{z,\lambda_{z}}-\hat{W}_{z}(s_{t-T},a_{t-T})\right)\nabla\hat{W}_{z}(s_{t-T},a_{t-T})$triangleright Update $\theta^{z}$ with TD (Eq.3) using $G^{z,\lambda_{z}}\forall z$.
              $\omega\leftarrow\omega+\alpha_{\omega}A^{\Delta}\nabla\log\pi_{\omega}(a_{t}|s_{t})$triangleright Update the policy parameters $\omega$ with PPO (Eq.4) for SARSA using $A^{\Delta}$
@@ 5 Analysis
Next, We examine our delta estimators through the lens of the bias-variance trade-off. In SARSA, bias emerges from dependence on current estimates, affecting updates over short time-scales, particularly with smaller discount factors, such as lower values of $\gamma$ in $W_{z}$. Updates utilizing smaller discount factors demonstrate reduced variance due to their reliance on current estimates; however, they exhibit increased bias as they fail to integrate substantial information regarding future rewards. In contrast, variance escalates for updates associated with extended time-scales, as future reward information adds further variability. In TD($\Delta$), action-value components $W_{z}$ with higher $\gamma_{z}$ are associated with greater long-term rewards; however, this results in increased variance stemming from the stochastic characteristics of reward sequences. We begin by demonstrating that our estimator is equivalent to the standard estimator $\hat{Q}_{\gamma}$ under specific conditions, as outlined in Theorem 1. This comparison elucidates the essential metrics of our estimator that may indicate potential advantages over the standard $\hat{Q}_{\gamma}$ estimator. Building on this result and previous research by Kearns and Singh[9], we modify these analyses for SARSA to investigate the effects of bias and variance, along with the TD($\Delta$) decomposition framework, in relation to action-value functions. Our objective is to extend the bias-variance error bound framework to action-value settings using TD($\Delta$)(i.e., Theorem 4), thereby offering a better understanding of how these quantities can be balanced to obtain optimal results[8].
@@ 5.1 SARSA Equivalency Configurations and Enhancements
$\hat{Q}(s,a)_{\gamma}:=\langle\theta^{\gamma},\phi(s,a)\rangle\quad\text{and}\quad\hat{W}_{z}(s,a):=\langle\theta^{z},\phi(s,a)\rangle,\forall z$
where $\theta$ and $\theta^{z}$ represent weight vectors in $\mathbb{R}^{d}$ and the function $\phi:S\times A\rightarrow\mathbb{R}^{d}$ represents a mapping from a state-action pair to a specified d-dimensional feature space. The weight vector $\theta$ for SARSA is updated according to the TD($\lambda$) learning rule in the following manner:
$\theta_{t+1}^{\gamma}=\theta_{t}^{\gamma}+\alpha\bigg{(}G_{t}^{\gamma,\lambda}-\hat{Q}_{\gamma}(s,a)\bigg{)}\phi(s_{t},a_{t}),$ (21)
where $G_{t}^{\gamma,\lambda}$ represents the TD($\lambda$) return defined in Eq.17. Likewise, the delta estimator approach TD($\lambda_{z},\Delta$) is used to update each $\hat{W}_{z}$:
$\theta_{t+1}^{z}=\theta_{t}^{z}+\alpha\bigg{(}G_{t}^{Z,\lambda_{z}}-\hat{W}_{z}(s_{t},a_{t})\bigg{)}\phi(s_{t},a_{t}),$ (22)
where $G_{t}^{Z,\lambda_{z}}$ is specified identically to the TD($\Delta$) return defined in Eq.19, modified for the particular action-value function. In these equations, $\alpha$ and $\{\alpha_{z}\}_{z}$ represent positive learning rates. The following theorem demonstrates the equivalence of two SARSA version algorithms in this context. The subsequent theorems resemble those presented by Romoff et al. [8]; however, we offer a proof concerning the action-value function in relation to SARSA, whereas they established the proof for the value function.
@@
If $\alpha_{z}=\alpha$, $\lambda_{z}\gamma_{z}=\lambda_{\gamma}$, $\forall z$, and if we choose the initial conditions in such a way that $\sum_{z=0}^{Z}\theta_{0}^{\gamma}=\theta_{0}^{\gamma}$, then the iterates produced by TD($\lambda$) (Eq. 21) and TD($\lambda,\Delta$) (Eq. 22) with linear function approximation satisfy the following conditions:
$\Sigma_{z=0}^{Z}\theta_{t}^{z}=\theta_{t}^{\gamma}$ (23)
@@
Equivalence is attained when $\lambda_{z}\gamma_{z}=\lambda_{\gamma},\forall z$. When $\lambda$ approaches 1 and $\gamma_{z}$ is less than $\gamma$, this condition suggests that $\lambda_{z}=\lambda_{\gamma}/\gamma_{z}$ could potentially surpass one, resulting in a risk of divergence in TD($\lambda_{z}$). The subsequent theorem demonstrates that the TD($\lambda$) operator, as defined in Eq.18, qualifies as a contraction mapping for the range $1\leq\lambda<\frac{1+\gamma}{2\gamma}$, thereby confirming that $\lambda_{\gamma}<1$[8].
@@
$\forall\lambda\in\left[0,\frac{1+\gamma}{2\gamma}\right]$, the operator $T_{\lambda}$ is defined by the Eq. $T_{\lambda}Q=Q+(I-\lambda\gamma P)^{-1}(TQ-Q)$ $,\quad\forall Q\in\mathbb{R}^{|S|\times|A|}$, is well-defined. Furthermore, $T_{\lambda}Q$ constitutes a contraction in relation to the max norm, with its contraction coefficient expressed as $\frac{\gamma}{|1-\lambda\gamma|}$.
@@
The analysis presented in Theorem 1 can be adapted to a new context in which $k_{z}$-step TD ($\Delta$) is employed for each $W_{z}$ component, rather than TD($\lambda$, $\Delta$), a variant of TD learning. Theorem 1 demonstrates that with linear function approximation, standard multi-step TD and multi-step TD($\Delta$) can be equivalent when the number of steps ($k_{z}$) is consistent across all time scales (i.e., $k_{z}=k,\forall z$).
@@
In addition to the learning rate, the approach of decomposing the action-value function into various components (specifically, $W_{z}$ components) presents advantages that traditional TD learning, including non-delta estimators or higher $W_{z}$ components, lacks. The ability to utilize various k-step returns (or $\lambda$-return) across different time scales provides enhanced control and adaptability in the learning process. If $k_{z}<k_{z+1},\forall z\quad(or\quad\gamma_{z}\lambda_{z}<\gamma_{z+1}\lambda_{z+1},\forall z)$, this approach may decrease variance while introducing bias, as shorter time scales, which can be learned more rapidly, may not align perfectly with longer time scales in terms of the values they estimate.
5.2 Evaluation for Minimizing $k_{z}$ Values in SARSA via Phased Updates
To show that how our approach differ from the single estimator case, let’s follow the tabular phased version of k-step TD presented by kearns & Singh[9]. In SARSA, we aim to estimate the value of a state-action pair (s, a) start from each state $s\in\mathcal{S}$ and with action $a\in\mathcal{A}$, then n trajectories generated $\bigg{\{}S_{0}^{(j)}=s,a_{0},\gamma_{0},...,S_{k}^{(j)},a_{k}^{(j)},\gamma_{k}^{(j)},S_{k+1}^{(j)},...\bigg{\}}_{1\leq j\leq n}$ by following the policy $\pi$ and averaging over trajectories. Let $Q_{\gamma,t}(s,a)$ denote the Q-value estimate at phase t for the discount factor $\gamma$. For each iteration t, also called phase t, to get the phase-based estimate $Q_{\gamma,t}(s,a)$ for (s, a), we average over the trajectories, resulting in:
$\hat{Q}_{\gamma,t}(s,a)=\frac{1}{n}\sum_{j=1}^{n}\bigg{(}\sum_{i=0}^{k-1}\gamma^{i}r_{i}^{(j)}+\gamma^{k}\hat{Q}_{\gamma,t-1}(s_{k}^{(j)},a_{k}^{(j)})\bigg{)}$ (24)
Theorem 3 below from the theorems of kearns & Singh[9] but they show it for state value function while we show an upper bound on the error for action-value function estimation defined by $\Delta_{t}^{\hat{Q}_{\gamma}}=max_{s,a}\bigg{\{}\bigg{|}\hat{Q}_{\gamma,t}(s,a)-Q_{\gamma}(s,a)\bigg{|}\bigg{\}}$

### v2

@@ 4.4 SARSA TD($\lambda,\Delta$)
Eq.20 introduces the $\lambda$-return[23, 1], which combines rewards over multiple steps to create a target for TD($\lambda$) updates. The $\lambda$-return $G_{n}^{\eta,\lambda}$ is defined as:
$G_{n}^{\eta,\lambda}(s_{n},a_{n})=\hat{Q}_{\eta}(s_{n},a_{n})+\sum_{k=0}^{\infty}(\lambda\eta)^{k}\delta_{n+k}^{\eta}$ (20)
@@
$\mathrm{T}_{\lambda}Q(s_{n},a_{n})=Q(s_{n},a_{n})+(I-\lambda\eta P)^{-1}(\mathrm{T}Q(s_{n},a_{n})-Q(s_{n},a_{n}))$ (21)
Here,P is the transition matrix for state-action pairs under the policy $\pi$, and Q is the action-value function. Similarly, Eq.22 defines the $\lambda$-return specific to delta estimators in TD($\lambda,\Delta$), denoted as $G_{n}^{m,\lambda_{m}}$ for each delta estimator $D_{m}$:
$G_{n}^{m,\lambda_{m}}:=\hat{D}_{m}(s_{n},a_{n})+\sum_{k=0}^{\infty}(\lambda_{m}\eta_{m})^{k}\delta_{n+k}^{m}$ (22)
where $\delta_{n}^{0}:=\delta_{n}^{\eta_{0}}$ and $\delta_{n}^{m}:=(\eta_{m}-\eta_{m-1})\hat{Q}_{\eta_{m-1}}(s_{n+1},a_{n+1})+\eta_{m}\hat{D}_{m}(s_{n+1},a_{n+1})-\hat{D}_{m}(s_{n+1},a_{n+1})$ are the TD-errors.
@@ 4.5 SARSA TD($\lambda,\Delta$) with Generalized Advantage Estimation (GAE)
Generalized Advantage Estimation (GAE) [28] seeks to compute the advantage function by aggregating multi-step TD errors. To apply this approach to delta estimators $D_{m}(s_{n},a_{n})$, we compute advantage estimates $A^{\Delta}(s_{n},a_{n})$ for each time-scale. Each advantage estimate $A^{\Delta}(s_{n},a_{n})$ employs a multi-step TD error pertinent to the delta estimator $D_{n}$, expressed as follows:
$A^{\Delta}(s_{n},a_{n})=\sum_{k=0}^{T-1}(\lambda_{m}\eta_{m})^{k}\delta_{n+k}^{\Delta}$ (23)
where $\delta_{n+k}^{\Delta}:=r_{n}+\eta_{m}\sum_{m=0}^{M}\hat{D}_{m}(s_{n+1},a_{n+1})-\sum_{m=0}^{M}\hat{D}_{m}(s_{n},a_{n})$.
We utilize $\eta_{m}$ as the discount factor and the sum of all D estimators as a surrogate for $Q_{\eta m}$. This objective is applicable to PPO by implementing the policy update from Eq.7 and replacing A with $A^{\Delta}$. Additionally, to train each $D_{m}$, we utilize a truncated form of their corresponding $\lambda$-return, as outlined in Eq.22. For more information, see Algorithm 2.
@@
Inputs: Pick out the discount factors ($\eta_{0},\eta_{1},...,\eta_{M}$), bootstrapping steps ($k_{0},k_{1},...,k_{M}$), and learning rates ($\alpha_{0},\alpha_{1},...,\alpha_{M}$).
Set the initial value of $D_{m}(s,a)=0$ for all states s, actions a, and scales m.
Initialize policy $\nu$, and values $\theta^{m}\forall m$
@@
     for n = 0, 1, 2,… do $\triangleright$ Loop for each time step
         Take action $a_{n}$, observe reward $r_{n}$ and next state $s_{n+1}$.
         Select next action $a_{n+1}$ based on a policy.
         for m = 0, 1, …, M do
              if $n\geq T$ then
                  $G^{m,\lambda_{m}}\leftarrow\hat{D}_{m}(s_{n-T},a_{n-T})+\sum_{k=0}^{T-1}(\lambda_{m}\eta_{m})^{k}\delta_{n-T+k}^{m}\forall m$triangleright Computing multi-step return $G^{m,\lambda_{m}}$ and TD-error $\delta_{n-T+k}^{m}$ using Eq.22.
@@
         for m = 0, 1, 2,…, M do
              $\hat{D}_{m}(s_{n-T},a_{n-T})\leftarrow\hat{D}_{m}(s_{n-T},a_{n-T})+\alpha_{m}\left(G^{m,\lambda_{m}}-\hat{D}_{m}(s_{n-T},a_{n-T})\right)$
              $A^{\Delta}=\sum_{k=0}^{T-1}(\lambda_{m}\eta_{m})^{k}\delta_{n-T+k}^{\Delta}$triangleright Where $A^{\Delta}$ and $\delta_{n-T+k}^{\Delta}$ are computed using Eq.23.
              $\theta^{m}\leftarrow\theta^{m}+\alpha_{m}\left(G^{m,\lambda_{m}}-\hat{D}_{m}(s_{n-T},a_{n-T})\right)\nabla\hat{D}_{m}(s_{n-T},a_{n-T})$triangleright Update $\theta^{m}$ with TD (Eq.3) using $G^{m,\lambda_{m}}\forall m$.
              $\mathcal{L}(\nu)=\mathbb{E}\bigg{[}min\bigg{(}\rho(\nu)A^{\Delta}(s,a),clip(\rho(\nu),1-\epsilon,1+\epsilon)A^{\Delta}(s,a)\bigg{)}\bigg{]}$triangleright from Eq.7
              $\nu\leftarrow\nu+\alpha_{\nu}\nabla_{\nu}\mathcal{L}(\nu)$triangleright Update the policy parameters $\nu$ with PPO (Eq.7) for SARSA using $A^{\Delta}$
@@ 5 Analysis
Next, We examine our delta estimators through the lens of the bias-variance trade-off. In SARSA, bias emerges from dependence on current estimates, affecting updates over short time-scales, particularly with smaller discount factors, such as lower values of $\eta$ in $D_{m}$. Updates utilizing smaller discount factors demonstrate reduced variance due to their reliance on current estimates; however, they exhibit increased bias as they fail to integrate substantial information regarding future rewards. In contrast, variance escalates for updates associated with extended time-scales, as future reward information adds further variability. In TD($\Delta$), action-value components $D_{m}$ with higher $\eta_{m}$ are associated with greater long-term rewards; however, this results in increased variance stemming from the stochastic characteristics of reward sequences. We begin by demonstrating that our estimator is equivalent to the standard estimator $\hat{Q}_{\eta}$ under specific conditions, as outlined in Theorem 1. This comparison elucidates the essential metrics of our estimator that may indicate potential advantages over the standard $\hat{Q}_{\eta}$ estimator. Building on this result and previous research by Kearns and Singh[9], we modify these analyses for SARSA to investigate the effects of bias and variance, along with the TD($\Delta$) decomposition framework, in relation to action-value functions. Our objective is to extend the bias-variance error bound framework to action-value settings using TD($\Delta$)(i.e., Theorem 4), thereby offering a better understanding of how these quantities can be balanced to obtain optimal results[8].
@@ 5.1 SARSA Equivalency Configurations and Enhancements
$\hat{Q}(s,a)_{\eta}:=\langle\theta^{\eta},\phi(s,a)\rangle\quad\text{and}\quad\hat{D}_{m}(s,a):=\langle\theta^{m},\phi(s,a)\rangle,\forall m$
where $\theta$ and $\theta^{m}$ represent weight vectors in $\mathbb{R}^{d}$ and the function $\phi:S\times A\rightarrow\mathbb{R}^{d}$ represents a mapping from a state-action pair to a specified d-dimensional feature space. The weight vector $\theta$ for SARSA is updated according to the TD($\lambda$) learning rule in the following manner:
$\theta_{n+1}^{\eta}=\theta_{n}^{\eta}+\alpha\bigg{(}G_{n}^{\eta,\lambda}-\hat{Q}_{\eta}(s,a)\bigg{)}\phi(s_{n},a_{n}),$ (24)
where $G_{n}^{\eta,\lambda}$ represents the TD($\lambda$) return defined in Eq.20. Likewise, the delta estimator approach TD($\lambda_{m},\Delta$) is used to update each $\hat{D}_{m}$:
$\theta_{n+1}^{m}=\theta_{n}^{m}+\alpha\bigg{(}G_{n}^{M,\lambda_{m}}-\hat{D}_{m}(s_{n},a_{n})\bigg{)}\phi(s_{n},a_{n}),$ (25)
where $G_{n}^{M,\lambda_{m}}$ is specified identically to the TD($\Delta$) return defined in Eq.22, modified for the particular action-value function. In these equations, $\alpha$ and $\{\alpha_{m}\}_{m}$ represent positive learning rates. The following theorem demonstrates the equivalence of two SARSA version algorithms in this context. The subsequent theorems resemble those presented by Romoff et al. [8]; however, we offer a proof concerning the action-value function in relation to SARSA, whereas they established the proof for the value function.
@@
Inspired by Romoff et al.[8], if $\alpha_{m}=\alpha$, $\lambda_{m}\eta_{m}=\lambda_{\eta}$, $\forall m$, and if we choose the initial conditions in such a way that $\sum_{m=0}^{M}\theta_{0}^{\eta}=\theta_{0}^{\eta}$, then the iterates produced by TD($\lambda$) (Eq. 24) and TD($\lambda,\Delta$) (Eq. 25) with linear function approximation satisfy the following conditions:
$\Sigma_{m=0}^{M}\theta_{n}^{m}=\theta_{n}^{\eta}$ (26)
@@
Equivalence is attained when $\lambda_{m}\eta_{m}=\lambda_{\eta},\forall m$. When $\lambda$ approaches 1 and $\eta_{m}$ is less than $\eta$, this condition suggests that $\lambda_{m}=\lambda_{\eta}/\eta_{m}$ could potentially surpass one, resulting in a risk of divergence in TD($\lambda_{m}$). The subsequent theorem demonstrates that the TD($\lambda$) operator, as defined in Eq.21, qualifies as a contraction mapping for the range $1\leq\lambda<\frac{1+\eta}{2\eta}$, thereby confirming that $\lambda_{\eta}<1$[8].
@@
Inspired by Romoff et al.[8], $\forall\lambda\in\left[0,\frac{1+\eta}{2\eta}\right]$, the operator $T_{\lambda}$ is defined by the Eq. $T_{\lambda}Q=Q+(I-\lambda\eta P)^{-1}(TQ-Q)$ $,\quad\forall Q\in\mathbb{R}^{|S|\times|A|}$, is well-defined. Furthermore, $T_{\lambda}Q$ constitutes a contraction in relation to the max norm, with its contraction coefficient expressed as $\frac{\eta}{|1-\lambda\eta|}$.
@@
The analysis presented in Theorem 1 can be adapted to a new context in which $k_{m}$-step TD ($\Delta$) is employed for each $D_{m}$ component, rather than TD($\lambda$, $\Delta$), a variant of TD learning. Theorem 1 demonstrates that with linear function approximation, standard multi-step TD and multi-step TD($\Delta$) can be equivalent when the number of steps ($k_{m}$) is consistent across all time scales (i.e., $k_{m}=k,\forall m$).
@@
Besides the learning rate, the method of dissecting the action-value function into distinct components (particularly, $D_{m}$ components) offers benefits that conventional TD learning, including non-delta estimators or elevated $D_{m}$ components, does not possess. The capacity to employ diverse k-step returns (or $\lambda$-return) across many time-scales offers improved control and flexibility in the learning process. If $k_{m}<k_{m+1},\forall m\quad(or\quad\eta_{m}\lambda_{m}<\eta_{m+1}\lambda_{m+1},\forall m)$, this approach may decrease variance while introducing bias, as shorter time-scales, which can be learned more rapidly, may not align perfectly with longer time scales in terms of the values they estimate.
5.2 Evaluation for Minimizing $k_{m}$ Values in SARSA via Phased Updates
To show that how our approach differ from the single estimator case, let’s follow the tabular phased version of k-step TD presented by kearns & Singh[9]. In SARSA, we aim to estimate the value of a state-action pair (s, a) start from each state $s\in\mathcal{S}$ and with action $a\in\mathcal{A}$, then n trajectories generated $\bigg{\{}S_{0}^{(x)}=s,a_{0},\eta_{0},...,S_{k}^{(x)},a_{k}^{(x)},\eta_{k}^{(x)},S_{k+1}^{(x)},...\bigg{\}}_{1\leq x\leq n}$ by following the policy $\pi$ and averaging over trajectories. Let $Q_{\eta,n}(s,a)$ denote the Q-value estimate at phase n for the discount factor $\eta$. For each iteration n, also called phase n, to get the phase-based estimate $Q_{\eta,n}(s,a)$ for (s, a), we average over the trajectories, resulting in:
$\hat{Q}_{\eta,n}(s,a)=\frac{1}{n}\sum_{x=1}^{n}\bigg{(}\sum_{i=0}^{k-1}\gamma^{i}r_{i}^{(x)}+\eta^{k}\hat{Q}_{\eta,n-1}(s_{k}^{(x)},a_{k}^{(x)})\bigg{)}$ (27)
Theorem 3 below from the theorems of kearns & Singh[9] but they show it for state value function while we show an upper bound on the error for action-value function estimation defined by $\Delta_{n}^{\hat{Q}_{\eta}}=max_{s,a}\bigg{\{}\bigg{|}\hat{Q}_{\eta,n}(s,a)-Q_{\eta}(s,a)\bigg{|}\bigg{\}}$